OpenStudy (anonymous):

What is 0.019 written as a percent? A. 0.00019% B. 0.019% C. 0.19% D. 1.9%

OpenStudy (anonymous):

move the decimal two places to the right 1.9%

OpenStudy (anonymous):

D

OpenStudy (anonymous):

0.019 times 100 then you get 1.9 This is how it works

OpenStudy (anonymous):

thanks

OpenStudy (whpalmer4):

A percentage \(x\%\) represents \(\dfrac{x}{100}\). We have \[0.019 = \frac{x}{100}\]Multiply both sides by 100\[0.019*100 = \frac{x}{100}*{100}\]\[1.9=x\]So the percentage is \(1.9\%\).